### v1

# STI-Bench: Are MLLMs Ready for Precise Spatial-Temporal World Understanding?

Abstract
The use of Multimodal Large Language Models (MLLMs) as an end-to-end solution for Embodied AI and Autonomous Driving has become a prevailing trend. While MLLMs have been extensively studied for visual semantic understanding tasks, their ability to perform precise and quantitative spatial-temporal understanding in real-world applications remains largely unexamined, leading to uncertain prospects. To evaluate models’ Spatial-Temporal Intelligence, we introduce STI-Bench, a benchmark designed to evaluate MLLMs’ spatial-temporal understanding through challenging tasks such as estimating and predicting the appearance, pose, displacement, and motion of objects. Our benchmark encompasses a wide range of robot and vehicle operations across desktop, indoor, and outdoor scenarios. The extensive experiments reveals that the state-of-the-art MLLMs still struggle in real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis.
Figure 1: We evaluate state-of-the-art MLLMs on STI-Bench for precise and quantitative spatial-temporal understanding using video inputs. Results indicate the significant challenge in all tasks.
Figure 2: Overview of STI-Bench. We selected the most representative videos from each dataset scene and provided a few simple questions for demonstration.
1 Introduction
The rapid development of Multimodal Large Language Models (MLLMs) [1, 12, 4, 34, 35, 36, 26, 32, 46, 41] has propelled them to the research forefront as a versatile tool to deal with numerous vision and multimodal tasks. Impressive performances have been achieved by MLLMs in general Visual Question Answering tasks [3], which mainly focus on the 2D visual perception and semantic question answering [18, 24, 37, 38, 19, 48, 20].
Beyond 2D visual perception, it has become a prevailing trend to employ MLLMs as an end-to-end solution for Embodied AI [22, 29, 25, 43, 15, 9, 8, 7] and Autonomous Driving [40, 44, 21, 39]. Such tasks require MLLMs to understand the 3D space and time, and then predict optimal manipulation strategies for robotic and vehicular systems. Although many explorations have been conducted, the question remains: Are MLLMs ready for precise spatial-temporal world understanding?
To answer this question, we propose a Spatial-Temporal Intelligence Benchmark (STI-Bench), designed to evaluate MLLMs’ spatial-temporal world understanding capability. We evaluate MLLMs using single video or multiple images as input instead of 3D point clouds. The main reasons are: 1) the majority of state-of-the-art models, e.g., GPT-4o [30] and Gemini [35], can accept images or video as input rather than 3D point clouds; 2) Videos are more frequently used in human’s daily life and they usually contain sufficient information to infer the spatial-temporal environment. STI-Bench contains 300 videos and more than 2,000 QA pairs, covering three major scenarios: Desktop, Indoor, and Outdoor. The videos are sourced from Omni6DPose [47], ScanNet [16] and Waymo [33] respectively, thus encompassing a broad spectrum of real-world environments. As illustrated in Figure 2, we design eight distinct tasks to evaluate models’ ability of static spatial measurement and grounding, and dynamic tasks including speed, acceleration and trajectory estimation.
Through extensive experiments as illustrated in Figure 1, we observe that even the most advanced MLLMs struggle with real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Our error analysis reveals three fundamental limitations: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal grounding and integration.
These insights highlight the significant challenges MLLMs face in precisely understanding spatial-temporal information from videos. We believe STI-Bench will serve as an important touchstone that guides the community to distinguish and develop better MLLMs for Embodied AI, Autonomous Driving tasks and beyond.
In summary, our main contributions include:
  • We present STI-Bench, comprising over 300 videos and more than 2,000 tailored questions across desktop, indoor, and outdoor scenarios, providing a systematic quantitative assessment of MLLMs’ spatial-temporal understanding capabilities.
  • We conduct an in-depth study of state-of-the-art video-based MLLMs on STI-Bench, identify key error patterns in spatial-temporal reasoning, and provide empirical insights that can help the community develop more reliable MLLMs for embodied applications.
Benchmark QA Pairs Data Env. Scene View Evaluation Spatial-Temporal D I O Ego Allo. Num. Desc. Dist. Dir. Vel. Traj. SAT [31] 218k I S VSI-Bench [42] 5,156 V R EmbSpatial-Bench [17] 3,640 I R EmbodiedAgentInterface [25] 448 - S - - EmbodiedEval [15] 328 I/V S - - EmbodiedBench [43] 1,128 I S - - WorldSense [6] 3,172 V R MLVU [48] 3,102 V R Video-MMMU [20] 300 V S STI-Bench 2,064 V R
Table 1: Comparison of STI-Bench with existing benchmarks. Data represents the source of our QA data, where V stands for Video and I stands for Image. Env. indicates the environment in which the data is generated, where S represents Simulation and R represents Real. The two columns under View indicate whether the dataset includes Ego-centric and Allocentric perspectives. The two columns under Evaluation specify whether the ground truth is presented in numerical or textual form. The four columns under Spatial-Temporal indicate whether the benchmark evaluates spatial distance, direction (with angular precision), velocity, or a precise and comprehensive trajectory description.
2 Related Work
2.1 Multimodal Large Language Models
Multimodal large language models (MLLMs) have achieved groundbreaking performance in visual understanding [1, 4, 12, 35], leveraging large language models (LLMs) [41, 36, 34] and visual encoders. Beyond image-based MLLMs, recent advancements have extended multimodal learning to video understanding. Classical works include models like VideoChat[23], which enable interactive video-based dialogue by integrating multimodal understanding. Subsequent models like Subsequent models like Video-LLaVA[27] enhance visual-language alignment through large-scale vision-language pretraining and fine-tuned adaptation, extending LLaVA[28]’s capability to process video inputs effectively. Recent works, Qwen2.5-VL [41] excels in long-video understanding and temporal localization by incorporating absolute temporal encoding, enabling the model to capture relationships among video frames more effectively. Additionally, its advancement in dynamic resolution modeling allows for seamless adaptation to videos with varying sampling rates, enhancing its versatility in processing diverse video inputs.
2.2 Spatial Understanding with MLLMs
Video MLLMs have attached great importance on semantic understanding. However, spatial understanding has always been a significant challenge, inspiring recent contribution [9, 11, 14]. This progress represents a significant step toward developing world models and embodied agents. Recent advancements in embodied intelligence have explored integrating large-scale MLLMs into robotic control, enabling better generalization and semantic reasoning. RT-2 [8] introduces a vision-language-action framework that transfers web-scale knowledge to robotic control by representing actions as tokens alongside visual and language data, allowing robots to generalize to novel objects and infer multi-step reasoning tasks. Building on this idea, GR-2 [10] extends generalist robot control across diverse embodiments using a Transformer-based architecture trained on a wide range of robotic tasks, demonstrating adaptability across different platforms. Further refining this approach, $\pi_{0}$ [7] incorporates a flow-matching mechanism to generate continuous, precise action trajectories, enhancing fine-grained manipulation skills. By integrating pretrained MLLMs with an independent action module, $\pi_{0}$ achieves zero-shot task execution and flexible adaptation through fine-tuning. Together, these models highlight the potential of leveraging large-scale learning for robotic control, pushing the boundaries of generalization, task adaptability, and multi-modal reasoning in embodied AI.
2.3 Video Benchmarks for MLLM
Recently, multiple benchmarks [38, 37, 48, 19] have emerged for comprehensively evaluating MLLMs’ ability of (long) video understanding, especially about visual perception and semantic reasoning in the form of Video Question Answering. LongVideoBench [38] and LVBench [37] focus on the understanding of long videos. Recent published benchmarks like Video-MME [19] and MMBench-Video [18] comprehensively evaluates MLLMs across various video-related tasks. Existing benchmarks primarily focus on high-level semantic understanding, such as entity recognition and event understanding. In addition, they are largely confined to a temporal extension of 2D image understanding, lacking precise 3D spatial and temporal reasoning of physical quantities. Recent works such as VSI-Bench [42], have shed light on a deeper understanding of the natural world by introducing visual-spatial intelligence tasks for MLLMs, where models are required to provide numerical answers in certain scenarios. However, as illustrated in Table 1, the limited inclusion of scenes and spatial-temporal tasks restricts their ability to capture the complexities of the real physical world. In contrast, STI-Bench comprehensively evaluate models’ ability of precise spatial-temporal understanding in tasks of static spatial measurement and physically motion understanding in Desktop, Indoor and Outdoor scenarios.
Figure 3: Benchmark Statistics. Top: Video length distribution across different categories and datasets. Bottom: The number of questions contributed by each dataset for evaluating different capabilities.
3 STI-Bench
In this section, we present the detailed design and construction of STI-Bench. The construction pipeline is depicted in Figure 4.
Figure 4: Benchmark curation pipeline. The pipeline first aggregates multi-scene RGB datasets that contain 3D bounding box annotations, camera parameters, and point cloud data, which serve as the basis for computing ground truth. From these datasets, we extract numerical ground truth such as distance and velocity, as well as textual descriptions of trajectories and directions. Subsequently, we leverage GPT to assist in generating QA pairs and design a website for rigorous human verification and filtering.
3.1 Task Definition
We propose eight tasks in total, each one systematically examining a distinct aspect of MLLMs’ spatial-temporal understanding. We divide these tasks into two main categories: Static Understanding and Dynamic Understanding.
Static Understanding
  1. a.
    Dimensional Measurement. Concerns estimates of an object’s geometric size, such as length, width, and height, as well as the distance between objects or between the camera and an object. This requires the ability to transform 2D pixel observations into physical world measurements and accurately perceive depth from monocular inputs.
    Example: ”What is the height of this box?” or ”How close is the camera to the table?”
  2. b.
    Spatial Relation. Focuses on identifying spatial relationships among objects or between the camera and an object, including front and back, left and right, up and below. This task tests models’ ability to understand relative positioning across different reference frames and maintain spatial relationship judgment consistency across varying viewpoints.
    Example: ”Is the chair on the left or right side of the table?” or ”What is the position of the red bag relative to the fur sofa?”
  3. c.
    3D Video Grounding. Given a semantic description such as ”the red backpack on the brown sofa,” the goal is to retrieve the object’s 3D bounding box in the camera coordinate system at a specific point in the video. This requires seamlessly aligning linguistic descriptions with visual features and accurately parameterizing 3D positional information.
    Example: ”Locate the 3D bounding box of the red suitcase near the bed.”
Dynamic Understanding
  1. d.
    Displacement and Path Length. Focuses on how far an object or the camera travels between two given time points. This requires tracking consistent reference points across frames and integrating motion information from discrete frames into continuous paths.
    Example: ”How far has the car traveled from 1s to 18s?”
  2. e.
    Speed and Acceleration. Investigates motion parameters by integrating spatial displacement with time intervals. This tests models’ ability to compute spatial derivatives with respect to time and maintain scale consistency across varying distances and perspectives.
    Example: ”What is the average speed of the camera?” or ”How quickly is the ball accelerating?”
  3. f.
    Ego-Centric Orientation. Examines how the camera’s azimuth orientation, parallel to the ground plane, changes over the duration of the video. This requires understanding rotation representations and utilizing fixed scene elements as angular reference points.
    Example: ”How many degrees does the camera’s horizontal orientation shift from the start of the video to its end?”
  4. g.
    Trajectory Description. Describes or infers the camera’s or an object’s motion path throughout the entire video, potentially involving multiple segments of travel and turns. This tests the ability to segment complex trajectories into meaningful components and abstract spatial motion patterns into concise language descriptions.
    Example: ”Summarize the camera trajectory, including distances moved and turns made.”
  5. h.
    Pose Estimation. Given the camera’s initial 3D pose, including position and orientation, estimates its pose at a specified point in the video using only the observed RGB data. This requires visual odometry capabilities and the ability to manage cumulative error in long sequences.
    Example: ”Given the initial pose of the camera, what is the camera’s pose at the requested time?”
Each of these tasks presents unique challenges that collectively evaluate models’ comprehensive spatial-temporal intelligence across different scales, from millimeter-precision desktop manipulation to meter-scale indoor navigation and beyond. Success in these tasks requires not only fundamental 3D spatial reasoning but also physical common sense and the ability to integrate information across different modalities and reference frames over time.
3.2 Benchmark Construction
Data Collection.
To encompass a broad spectrum of real-world environments, STI-Bench covers three major scenarios: Desktop, Indoor, and Outdoor. Accordingly, we draw from three publicly available datasets—Omni6DPose[47] for desktop-scale 6D object pose estimation, ScanNet[16] for indoor 3D scene reconstruction, and Waymo [33] for autonomous driving. These datasets provide frame-by-frame camera intrinsic and extrinsic parameters, as well as point clouds for each object, which we map to two-dimensional bounding boxes in each frame.
Automatic QA Pair Generation.
We used MLLMs to produce detailed semantic descriptions for each object, such as “A beige minivan with a roof rack,” “A refrigerator with emoji magnets, photos, and a to-do list,” or “A red backpack on a brown leather sofa.” Next, leveraging the frame-by-frame annotations, we computed the ground-truth information required for each task. We then provided the ground-truth data, object descriptions, and task-specific QA requirements to MLLMs to generate a diverse set of questions and challenging answer options.
Human Quality Control.
During QA pair generation, several issues arose:
  1. 1.
    LLM-generated descriptions could be inaccurate or fail to uniquely identify the target object.
  2. 2.
    Some questions and options remained unreasonable or incorrect, even with detailed guidelines.
  3. 3.
    In certain cases, the video alone did not provide sufficient information. For example, the camera was occluded but lidar data were available.
To address these challenges, we developed a website for multiple rounds of manual filtering and sampling-based review, ensuring high-quality questions. We also randomly shuffled the answer options to enhance evaluation robustness. Ultimately, we curated more than 2,000 high-quality QA pairs from over 300 videos. Details are shown in Figure 3.
Fine-Grained Adjustment.
After generating and refining the QA pairs, we recognized that real-world applications differ significantly in terms of error tolerance. For instance, a desktop robotic arm may require millimeter-level precision, whereas autonomous driving can function effectively with meter-scale accuracy. To accommodate these varied needs, we applied a scaling factor to the numerical differences between correct answers and distractors, aligning them with the precision requirements of specific scenarios. Consequently, the smallest margin of error ranges from millimeters to centimeters in desktop settings, centimeters to decimeters indoors, and decimeters to meters outdoors. We also adopted a logarithmic sampling approach to avoid clustering most differences at the higher bounds of each range. This fine-grained adjustment preserves the semantic value of each question while maintaining suitable gradients across different precision levels, enabling more effective training and evaluation of MLLMs in diverse environments and industries.
  Methods Rank Avg. Static Understanding Dynamic Understanding
Dim. Meas. Disp. & P.L. Speed & Acc. Spatial Relation Ego Orient. Traj. Desc. 3D Video Grounding Pose Est.
Proprietary Models (API)
GPT-4o[30] 6 28.3 20.1 20.6 26.9 48.0 22.2 35.9 21.8 42.8
Gemini-2.0-Flash[35] 4 35.6 30.8 25.1 32.6 48.0 20.5 48.7 28.1 59.1
Gemini-2.5-Pro[35] 2 40.5 32.9 30.8 34.0 53.4 55.7 42.9 36.2 52.5
Claude-3.7-Sonnet[2] 3 37.2 32.5 25.4 34.4 48.3 26.5 42.3 34.7 57.2
Open-source Models
Qwen2.5-VL-72B[5] 1 40.8 34.6 28.1 36.1 45.5 48.0 46.2 35.8 60.5
InternVL2.5-78B[13] 5 29.3 26.0 20.7 27.0 52.5 19.7 28.4 25.6 45.1
VideoLLaMA3-7B[45] 7 27.4 26.3 24.3 24.5 39.7 21.1 24.4 27.8 32.5
Table 2: Evaluation on Ourbench. Orange marks the best result, and Light Orange marks the second best.
Model Outdoor Indoor Desktop Overall
Claude-3.7-Sonnet 39.27 35.84 35.94 37.17
Gemini-2.0-Flash 39.97 34.22 30.07 35.61
Gemini-2.5-Pro 48.02 35.36 36.92 40.52
GPT-4o 36.95 21.27 26.16 28.25
InternVL2.5 33.39 28.85 23.79 29.34
Qwen2.5-VL 48.23 35.48 36.70 40.84
VideoLLaMA3-7B 26.99 28.79 25.12 27.37
Table 3: Evaluation results across different scenes.
Figure 5: Evaluation results across different scenes and tasks.
4 Experiments
4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess the performance of four proprietary models, GPT-4o[30], Gemini-2.0-Flash[35], Gemini-2.5-Pro[35], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13] and VideoLLaMA3-7B[45].
To ensure a consistent temporal sampling strategy across videos, we sample frames at 1 fps for both input and output. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields a 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth, without employing any additional external models or annotations for performance evaluation.
4.2 Main Results
As shown in Table 2, we present a comprehensive evaluation of various MLLMs on STI-Bench. Overall, Qwen2.5-VL-72B achieves the highest accuracy of 40.8% among all tested models, slightly outperforming Gemini-2.5-Pro (40.5%). While these results significantly exceed the random guess baseline (20%), they still highlight substantial room for improvement in spatial-temporal understanding.
When analyzing performance across different scene types, we observe consistent patterns. All models perform better in outdoor scenarios (Qwen2.5-VL: 48.2%, Gemini-2.5-Pro: 48.0%) compared to indoor (35.5%, 35.4%) and desktop environments (36.7%, 36.9%). This suggests that models might have been exposed to more outdoor video content during training, or that outdoor scenes often provide clearer visual cues for spatial relationships.
Task-specific performance reveals particularly challenging areas. Most models struggle significantly with Displacement & Path Length estimation (best: 31.6% by Qwen2.5-VL) and Dimensional Measurement (best: 34.6% by Qwen2.5-VL), both of which require precise quantitative spatial understanding. In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 60.5% by Qwen2.5-VL) and Spatial Relation tasks (best: 55.4% by Qwen2.5-VL).
Notably, open-source models like Qwen2.5-VL and InternVL2.5 demonstrate competitive performance compared to proprietary models, with Qwen2.5-VL even slightly outperforming Gemini-2.5-Pro on several metrics. However, smaller models like VideoLLaMA3-7B (27.4% overall) still lag significantly behind their larger counterparts.
It is important to emphasize that even the best-performing models achieve only about 40% accuracy on our benchmark, which, while twice the random guess baseline, remains far from the reliability required for real-world embodied AI or autonomous driving applications. These results indicate that current MLLMs, despite their impressive capabilities in general visual understanding, still struggle with precise spatial-temporal intelligence tasks central to embodied applications.
Figure 6: The simplified thought process examples of Gemini 2.5 Pro.
4.3 Experimental Analysis
Given that Gemini-2.5-Pro is a multi-modal reasoning model with detailed thinking processes and ranks second-best among all tested models (best among proprietary models), we select it as a representative for in-depth analysis. The simplified thought process examples is presented in Figure 6.
Our analysis of Gemini-2.5-Pro reveals several key characteristics of current state-of-the-art MLLMs’ spatial-temporal understanding capabilities. Overall, the model achieves 40.52% accuracy. Performance is notably stronger in outdoor scenarios (48.02%) compared to desktop (36.92%) and indoor environments (35.36%). This performance disparity suggests that the model’s training data likely emphasized outdoor scenes and larger-scale understanding, or that outdoor environments typically provide clearer visual cues for spatial reasoning.
When examining task-specific performance, we observe that Gemini-2.5-Pro demonstrates stronger capabilities in orientation and spatial relationship tasks. It achieves the highest accuracy in Ego-Centric Orientation (55.74%) and Spatial Relation tasks (53.42%), followed by Pose Estimation (52.53%). However, the model struggles significantly with tasks requiring precise quantitative estimation, particularly Displacement & Path Length (30.81%), Dimensional Measurement (32.87%), and Speed & Acceleration (34.04%).
By leveraging the model’s reasoning process and uniformly sampling approximately 200 error records across each task type and scenario, we categorize its errors into three representative patterns. Figure 7 shows the distribution of error categories.
Figure 7: Distribution of error categories in Gemini-2.5-Pro across our sampled error cases.
Inaccurate Spatial Quantification
The model faces significant challenges in accurately estimating static spatial properties and relationships from visual inputs. These difficulties manifest when estimating object dimensions (size), distances between objects or between camera and objects, and absolute 3D coordinates/bounding boxes at specific time points. These errors stem from a lack of clear visual size references, difficulty distinguishing between numerically close options, and the inherent challenges of inferring metric scale from 2D pixels and estimating depth with monocular cameras. Such limitations directly impact performance in dimensional measurement, spatial relation, and 3D video grounding tasks.
Flawed Temporal Dynamics Understanding
The model struggles to perceive, track, and interpret cross-frame information that changes over time, such as motion and its dynamics. This results in erroneous calculations or descriptions of displacement, path length, speed, acceleration, directional changes (ego-centric or object pose), and overall trajectory shapes. The model particularly struggles with relative motion (distinguishing object motion from camera motion), a problem exacerbated by sparse temporal sampling. These difficulties arise from challenges in integrating information across frames, lack of internal models for physics/kinematics, inability to separate ego-motion from object motion, and information loss due to sparse sampling. These issues manifest in tasks involving displacement and path length, speed and acceleration, ego-centric orientation, trajectory description, and pose estimation (as it changes over time).
Weak Cross-Modal Grounding and Integration
The model fails to properly connect textual queries/instructions with relevant spatial-temporal visual content, or to integrate provided non-visual data (such as initial poses) with visual information. This includes misinterpreting temporal constraints (like ”from 1s to 18s,” ”at the end,” ”the moment of last co-occurrence”), failing to correctly utilize provided initial conditions (e.g., initial camera pose in pose estimation tasks), and incorrectly associating structured data (coordinates, timestamps) with visual elements. These errors stem from deficiencies in parsing structured/natural language instructions and difficulty integrating information from different modalities (text prompts, initial state data, video frames) into a unified reasoning process. This affects all tasks that rely on specific instructions or initial data.
These error patterns highlight that, despite Gemini-2.5-Pro’s strong performance relative to other models, it still faces significant challenges in precise spatial-temporal understanding. It’s limitations in quantitative estimation and complex spatial-temporal reasoning indicate that current MLLMs remain far from achieving the reliability required for embodied AI or autonomous driving applications.
5 Conclusion
We introduced STI-Bench, a comprehensive benchmark to assess MLLMs’ spatial-temporal understanding through over 300 real-world videos and 2,000 QA pairs of robot desktop, indoor, and outdoor scenarios, which reveals significant limitations in current MLLMs’ spatial-temporal understanding capabilities, with even top-performing models achieving only 40-48% accuracy. Models particularly struggle with precise quantitative tasks like dimensional measurement. Our analysis identifies three key weaknesses: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal integration. These findings emphasize the substantial gap between current capabilities and the reliability needed for embodied AI and autonomous driving applications. STI-Bench provides a valuable framework for evaluating and improving MLLMs’ ability to understand the physical world—essential for developing the next generation of embodied intelligent systems.